OpenStudy (nightmarenight):

Help pls? c': Divide. (6a^4 + 2a^3) ÷ 2a @bibby ?

OpenStudy (bibby):

You can rewrite the fraction as \[\huge \frac{ 6a^4}{ 2a } +\frac{ 2a^3}{ 2a } \]

OpenStudy (bibby):

that's some interesting math you're working there Start with the left hand side. What is a^4/a You can rewrite it:\[\huge \frac{ a^4 }{ a }=\frac{ a*a*a*a }{ a }\]

OpenStudy (bibby):

Yeah. You can't add things with different exponents. it might help if you think of them as shapes or fruits. if a^4's are apples and a^3's are oranges You can only add apples+apples. oranges+oranges. not oranges + apples

OpenStudy (bibby):

anyhow \[\huge \frac{ a^4 }{ a }=\frac{ \cancel a*a*a*a }{ \cancel a }\]

OpenStudy (nightmarenight):

Okay, so now a^3? @bibby

OpenStudy (bibby):

that's not everything. We have the following: \[\huge \frac{ 6a^4}{ 2a } +\frac{ 2a^3}{ 2a }= \frac{ 6a^3}{ 2 } +\frac{ 2a^3}{ 2a }\] Let's finish the left side. What is 6/2?

OpenStudy (nightmarenight):

3 @bibby

OpenStudy (bibby):

:3 So\[\huge 3a^3 +\frac{ 2a^3}{ 2a }\] What can we do with the right side?

OpenStudy (nightmarenight):

divide? but could you put a 1 under 3a^3?

OpenStudy (bibby):

Do you want to add the two terms? You can't do that yet as they don't have common denominators.

OpenStudy (nightmarenight):

I'm not sure @bibby

OpenStudy (bibby):

OK take \[\huge\frac{ 2a^3}{ 2a }\]alone try simplifying that

OpenStudy (nightmarenight):

4a^4?

OpenStudy (nightmarenight):

wait

OpenStudy (bibby):

waiting

OpenStudy (nightmarenight):

1a^3???

OpenStudy (nightmarenight):

divide? I don't even know

OpenStudy (bibby):

yeah. you got it right. a^3

OpenStudy (bibby):

I'll write it out\[\huge\frac{ 2a^3}{ 2a }=\huge\frac{ \cancel 2a^{\cancel32}}{ \cancel2\cancel a }\]

OpenStudy (bibby):

actually you got it wrong, whoops

OpenStudy (nightmarenight):

^2?

OpenStudy (nightmarenight):

a^2 ***

OpenStudy (bibby):

Yeah. Recall the exponent division rule from earlier. a^3/a= a^(3-1) = a^2

OpenStudy (bibby):

\[\huge 3a^3 +{ a^2}\]and we're done

OpenStudy (bibby):

You can factor out a^2 though: a^2(3a + 1)

OpenStudy (nightmarenight):

Ohhhh. Okay, I wasn't getting the exponent division rule. Thank you!